OpenStudy (anonymous):

A 100 g ball moving to the right at 4.0 m/s collides head-on with a 200 g ball that is moving to the left at 3.0 m/s. If the collision is perfectly elastic, what are the speeds of each ball after the collision?

OpenStudy (anonymous):

\[(0.1kg)(4\frac{ m }{ s}) - (0.2kg)(3\frac{ m }{ s }) = -(0.1kg)v _{f1} + (0.2kg)v _{f2}\]

OpenStudy (anonymous):

\[-0.2Ns = -(0.1kg)v _{f1}+(0.2kg)v _{f2}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(0.1kg)(4\frac{ m }{ s })^{2}+\frac{ 1 }{ 2 }(0.2kg)(3\frac{ m }{ s })^{2} = K _{f}\]

OpenStudy (anonymous):

\[1.7J = \frac{ 1 }{ 2 }(0.1kg)v _{f1}^{2}+\frac{ 1 }{ 2 }(0.2kg)\frac{ (-0.2+0.1v _{f1}) }{ 0.2kg }\]

OpenStudy (anonymous):

Can someone verify if I am on the right track?